OpenStudy (megannicole51):

Find the center of mass of a cone of height 5cm and base diameter 10cm with constant density of gm/cm^3.

OpenStudy (anonymous):

v=1/3 pi r^2 h and d=m/v

OpenStudy (megannicole51):

to find the mass what is the integral going to? like what are a and b?

OpenStudy (megannicole51):

0 to 5?

OpenStudy (anonymous):

u mean center by center of mass?

OpenStudy (skullpatrol):